# A Framework for Safe Probabilistic Invariance Verification of Stochastic Dynamical Systems

### v1

@@ 5.2 Reachability Probability Characterization in [31]
$\begin{cases}v(\bm{x})\leq\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\leq 0&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ \mathcal{A}u(\bm{x})-v(\bm{x})\leq 0&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ -v(\bm{x})\leq 0,&\forall\bm{x}\in\partial{\mathcal{X}}\\ 1-v(\bm{x})\leq 0&\forall\bm{x}\in\mathcal{X}_{r}\end{cases},$ (24)
then $\mathbb{P}\big{(}\exists\tau\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(\tau,\bm{w})\in\mathcal{X}_{r}\wedge\forall t\in[0,\tau).\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}=\mathbb{P}\big{(}\exists\tau\in\mathbb{R}_{\geq 0}.\widetilde{\bm{X}}^{\bm{x}_{0}}(\tau,\bm{w})\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\leq\epsilon_{2}$ holds.
@@ 6 Safe Probabilistic Invariance Verification for Continuous-time Systems
In this section, we present optimizations to compute the the upper and lower bounds of the liveness probability of continuous-time systems. We first present an optimization for computing lower bounds of the liveness probability, which is adapted from the classical stochastic barrier certificate based on the Doob’s nonnegetive supermartingale inequality. We then present alternative optimizations for lower- and upper-bounding the liveness probability, which are adapted from the conditions in Propositions 5.23 and 5.24.
Similar to Theorem 5.22, all optimizations are based on a new stochastic process $\{\widehat{\bm{X}}^{\bm{x}_{0}}(t,\bm{w}),t\in\mathbb{R}_{\geq 0}\}$ for $\bm{x}_{0}\in\overline{\mathcal{X}}$, which is a stopped process corresponding to $\{\bm{X}^{\bm{x}_{0}}(t,\bm{w}),t\in[0,T^{\bm{x}_{0}}(\bm{w}))\}$ and the set $\mathcal{X}$, i.e.,
$\widehat{\bm{X}}^{\bm{x}_{0}}(t,\bm{w}))=\begin{cases}\bm{X}^{\bm{x}_{0}}(t,\bm{w})&\text{if}\ t<\tau^{x_{0}}(\bm{w})\\ \bm{X}^{\bm{x}_{0}}(\tau^{x_{0}}(\bm{w}),\bm{w})&\text{if}\ t\geq\tau^{x_{0}}(\bm{w})\end{cases},$ (25)
where
$\tau^{x_{0}}(\bm{w})=\inf\{\,t\mid\bm{X}^{\bm{x}_{0}}(t,\bm{w}))\in\partial\mathcal{X}\,\}$
is the first time of exit of $\bm{X}^{\bm{x}_{0}}(t,\bm{w})$ from the open set $\mathcal{X}$. Clearly, the set $\overline{\mathcal{X}}$ is an invariant set for $\widehat{\bm{X}}^{\bm{x}_{0}}(t,\bm{w})$ with $\bm{x}_{0}\in\overline{\mathcal{X}}$. Moreover, the infinitesimal generator corresponding to $\widehat{\bm{X}}^{\bm{x}_{0}}(t,\bm{w})$ (denoted $\widehat{\mathcal{A}}$) is identical to the one corresponding to $\bm{X}^{\bm{x}_{0}}(t,\bm{w})$ on the set $\mathcal{X}$, and is equal to zero on the set $\partial\mathcal{X}$, i.e.,
@@
6.1 Barrier Certificates Based Invariance Verification
This subsection formulates a sufficient condition for lower-bounding the liveness probability in Definition 5.21. This condition is the straightforward extension of the one in Theorem 15 in [17] and the one in Proposition III.8 in [27], which is built upon the well established Doob’s nonnegative supermartingale inequality [26]. This condition is formulated here primarily for convenient comparisons throughout the remainder.
@@ 6.1 Barrier Certificates Based Invariance Verification
$\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\leq 0&\forall\bm{x}\in\overline{\mathcal{X}}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\overline{\mathcal{X}}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\mathcal{X}_{unsafe}(=\partial\mathcal{X})\end{cases},$ (26)
then
$\mathbb{P}\big{(}\exists t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\partial\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\leq 1-\epsilon_{1}.$
Thereby, $\mathbb{P}\big{(}\forall t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}\geq 1-v(\bm{x}_{0})\geq\epsilon_{1}$
According to Proposition 6.25, an lower bound of the liveness probability can be computed via solving the following optimization:
$\begin{split}&\textbf{Op5}~{}~{}~{}~{}~{}~{}~{}\max_{v(\bm{x}),\epsilon_{1}}\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\leq 0&\forall\bm{x}\in\overline{\mathcal{X}}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\overline{\mathcal{X}}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\partial\mathcal{X}\\ \epsilon_{1}\geq 0\end{cases}.\end{split}$
To the best of our knowledge, there are no barrier-like conditions based on the Doob’s nonnegative supermartingale inequality, similar to the one stated in Theorem 16 of [3] for discrete-time systems, that have been developed in previous studies to examine the upper bounds of the reachability probability. Moreover, it is not possible to adapt the constraint $\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]-v(\bm{x})\leq-\delta,\forall\bm{x}\in\mathcal{X}$ from equation (10) to $\mathcal{A}v(\bm{x})\leq-\delta,\forall\bm{x}\in\mathcal{X}$ in order to compute upper bounds for the liveness probability. If we consider that $\mathcal{A}v(\bm{x})\leq-\delta,\forall\bm{x}\in\mathcal{X}$ is true, then $\widehat{\mathcal{A}}v(\bm{x})\leq-\delta,\forall\bm{x}\in\mathcal{X}$ will also hold, which implies $\widehat{\mathcal{A}}v(\bm{x})\leq-\delta,\forall\bm{x}\in\overline{\mathcal{X}}$. However, this contradicts the condition $\widehat{\mathcal{A}}v(\bm{x})=0,\forall\bm{x}\in\partial\mathcal{X}$.
@@ 6.2 Equations Relaxation Based Invariance Verification
In this subsection, we present the second set of optimizations for addressing the safe probabilistic invariance verification problem in Definition 5.21. We begin by introducing an equation that characterizes the liveness probability of satying the safe set $\mathcal{X}$ for all the time. This equation is adapted from (20), where the unsafe set $\partial\mathcal{X}$ is regarded as the target set $\mathcal{X}_{r}$ in (20). We then propose two sufficient conditions for certifying lower and upper bounds of the liveness probability by relaxing the derived equation.
Theorem 6.26.
@@
$\begin{cases}\widehat{\mathcal{A}}v(\bm{x})=0\\ v(\bm{x})=1_{\partial\mathcal{X}}(\bm{x})+\widehat{\mathcal{A}}u(\bm{x})\end{cases},$ (27)
@@
Thereby, $\mathbb{P}\big{(}\forall t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}\big{)}=1-v(\bm{x}_{0})$
Proof 6.27.
@@
Similar to the Section 4, two sufficient conditions can be obtained for certifying lower and upper bounds of the liveness probability via directly relaxing equation (27), respectively.
@@
$\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\leq 0&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq\mathcal{A}u(\bm{x})&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\partial{\mathcal{X}}\end{cases},$ (28)
then $\mathbb{P}\big{(}\exists t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\partial\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\leq v(\bm{x})\leq 1-\epsilon_{1}$. Hence, $\mathbb{P}\big{(}\forall t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\geq\epsilon_{1}$
Proof 6.29.
@@
Similarly, we can show that (26) and (28) are equivalent.
Lemma 6.30.
Constraints (26) and (28) are equivalent.
Proof 6.31.
The proof is similar to Lemma 4.15, which is shown in Appendix.
According to Proposition 6.28, a lower bound of the liveness probability can be computed via solving the following optimization:
$\begin{split}&\textbf{Op6}~{}~{}~{}~{}~{}~{}~{}\max_{v(\bm{x}),u(\bm{x}),\epsilon_{1}}\epsilon_{1}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\leq 0&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq\mathcal{A}u(\bm{x})&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\partial{\mathcal{X}}\\ \epsilon_{1}\geq 0\\ u(\bm{x})\text{~{}is bounded over~{}}\overline{\mathcal{X}}\end{cases}.\end{split}$
The sufficient condition for certifying upper bounds of the liveness probability is presented in Proposition 6.32.
Proposition 6.32.
$\begin{cases}v(\bm{x})\geq 1-\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq\mathcal{A}u(\bm{x})&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq 1&\forall\bm{x}\in\partial{\mathcal{X}}\end{cases},$ (29)
then $\mathbb{P}\big{(}\exists t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\partial\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\geq v(\bm{x})\geq 1-\epsilon_{2}$. Hence, $\mathbb{P}\big{(}\forall t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\mid\bm{x}_{0}\in\mathcal{X}_{0}\big{)}\leq\epsilon_{2}$.
Proof 6.33.
According to Proposition 6.32, an upper bound of the liveness probability can be computed via solving the following optimization:
$\begin{split}&\textbf{Op7}~{}~{}~{}~{}~{}~{}~{}\min_{v(\bm{x}),u(\bm{x}),\epsilon_{2}}\epsilon_{2}\\ &\text{s.t.~{}}\begin{cases}v(\bm{x})\geq 1-\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq\mathcal{A}u(\bm{x})&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq 1&\forall\bm{x}\in\partial{\mathcal{X}}\\ \epsilon_{2}\geq 0\\ u(\bm{x})\text{~{}is bounded over~{}}\overline{\mathcal{X}}\end{cases}.\end{split}$
@@ 7 Examples
In this section we demonstrate our theoretical developments on four examples. Since directly solving the problems Op0-Op8 is challenging, we relax them into semi-definite programs (SDP0-SDP8, see Appendix) using the sum of squares decomposition for multivariate polynomials, which allows for efficient solving. We solve the resulting semi-definite programs using the tool Mosek 10.1.21 [4]. Furthermore, to ensure numerical stability during the solution of the semi-definite programs, we impose a constraint on the coefficients of the unknown polynomials ($v(\bm{x}),w(\bm{x}),u(\bm{x}),p(\bm{x}),s_{i}(\bm{x}),i=0,\ldots,4$) in SDP0-SDP8 and restrict them to the interval $[-100,100]$.
In addition, in the experiments we employ the Monte Carlo method to approximate the upper and lower bounds of the liveness probability. Specifically, we take a large number of samples from the initial set and simulate the liveness probability of these samples over an extended time. The maximum probability among these samples is considered as the upper bound, while the minimum probability is considered as the lower bound. By using these bounds as ground truth values, we can assess the accuracy of the bounds computed via solving SDP0-SDP8.
Op3 and Op4
d 2 4 6 8 10 12 14 16 18 20 22 24 26
$\epsilon_{1}$ 0.3574 0.5890 0.6678 0.6895 0.6917 0.7281 0.7368 0.7549 0.7575 0.7597 0.7622 0.7630 0.7647
$\epsilon_{2}$ 1.0000 0.9844 0.9505 0.9489 0.9488 0.9474 0.9242 0.9143 0.8991 0.8991 0.8927 0.8804 0.8771
Op1
$\epsilon_{1}$ 0.3574 0.5890 0.6678 0.6895 0.6917 0.7281 0.7368 0.7549 0.7575 0.7597 0.7622 0.7630 0.7647
Op0
$\epsilon_{1}$ 0.3574 0.5890 0.5895 0.5902 0.5937 0.6192 0.6740 0.6863 0.6867 0.6887 0.7007 0.7259 0.7284

### v2

@@ 5.2 Reachability Probability Characterization in [32]
(24) $\begin{cases}v(\bm{x})\leq\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\leq 0&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ \mathcal{A}u(\bm{x})-v(\bm{x})\leq 0&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ -v(\bm{x})\leq 0,&\forall\bm{x}\in\partial{\mathcal{X}}\\ 1-v(\bm{x})\leq 0&\forall\bm{x}\in\mathcal{X}_{r}\end{cases},$
then
$\begin{split}&\mathbb{P}\big{(}\exists\tau\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(\tau,\bm{w})\in\mathcal{X}_{r}\wedge\forall t\in[0,\tau).\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\big{)}\\ &=\mathbb{P}\big{(}\exists\tau\in\mathbb{R}_{\geq 0}.\widetilde{\bm{X}}^{\bm{x}_{0}}(\tau,\bm{w})\in\mathcal{X}_{r}\big{)}\leq\epsilon_{2},\forall\bm{x}_{0}\in\mathcal{X}_{0}.\end{split}$
@@ 6 Safe Probabilistic Invariance Verification for Continuous-time Systems
In this section, we present optimizations to compute the the upper and lower bounds of the liveness probabilities of continuous-time systems. We first present an optimization for computing lower bounds of the liveness probabilities, which is adapted from the classical stochastic barrier certificate based on the Doob’s nonnegetive supermartingale inequality. We then present alternative optimizations for lower- and upper-bounding the liveness probabilities, which are adapted from the conditions in Corollary 5.5 and 5.6.
Similar to Proposition 5.4, all optimizations are based on a new stochastic process $\{\widehat{\bm{X}}^{\bm{x}_{0}}(t,\bm{w}),t\in\mathbb{R}_{\geq 0}\}$ for $\bm{x}_{0}\in\overline{\mathcal{X}}$, which is a stopped process corresponding to $\{\bm{X}^{\bm{x}_{0}}(t,\bm{w}),t\in[0,T^{\bm{x}_{0}}(\bm{w}))\}$ and the set $\mathcal{X}$, i.e.,
(25) $\widehat{\bm{X}}^{\bm{x}_{0}}(t,\bm{w}))=\begin{cases}\bm{X}^{\bm{x}_{0}}(t,\bm{w})&\text{if}\ t<\tau^{x_{0}}(\bm{w})\\ \bm{X}^{\bm{x}_{0}}(\tau^{x_{0}}(\bm{w}),\bm{w})&\text{if}\ t\geq\tau^{x_{0}}(\bm{w})\end{cases},$
where $\tau^{x_{0}}(\bm{w})=\inf\{\,t\mid\bm{X}^{\bm{x}_{0}}(t,\bm{w}))\in\partial\mathcal{X}\,\}$ is the first time of exit of $\bm{X}^{\bm{x}_{0}}(t,\bm{w})$ from the open set $\mathcal{X}$. Clearly, the set $\overline{\mathcal{X}}$ is an invariant set for $\widehat{\bm{X}}^{\bm{x}_{0}}(t,\bm{w})$ with $\bm{x}_{0}\in\overline{\mathcal{X}}$. Moreover, the infinitesimal generator corresponding to $\widehat{\bm{X}}^{\bm{x}_{0}}(t,\bm{w})$ (denoted $\widehat{\mathcal{A}}$) is identical to the one corresponding to $\bm{X}^{\bm{x}_{0}}(t,\bm{w})$ on the set $\mathcal{X}$, and is equal to zero on the set $\partial\mathcal{X}$, i.e.,
@@
6.1 Doob’s Nonnegative Supermartingale Inequality Based Invariance Verification
This subsection formulates a sufficient condition for lower-bounding the liveness probabilities in Definition 5.3. This condition is the straightforward extension of the one in Theorem 15 in [18] and the one in Proposition III.8 in [28], which is built upon the well established Doob’s nonnegative supermartingale inequality [27]. This condition is formulated here primarily for convenient comparisons throughout the remainder.
@@ 6.1 Doob’s Nonnegative Supermartingale Inequality Based Invariance Verification
(26) $\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\leq 0&\forall\bm{x}\in\overline{\mathcal{X}}\\ v(\bm{x})\geq 0&\forall\bm{x}\in\overline{\mathcal{X}}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\mathcal{X}_{unsafe}(=\partial\mathcal{X})\end{cases},$
then $\mathbb{P}\big{(}\exists t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\partial\mathcal{X}\big{)}\leq 1-\epsilon_{1}$ and thus $\mathbb{P}\big{(}\forall t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\big{)}\geq 1-v(\bm{x}_{0})\geq\epsilon_{1}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$.
According to Proposition 6.1, a lower bound of the liveness probabilities can be computed via solving Op5 (shown in Appendix B).
To the best of our knowledge, there are no barrier-like conditions based on the Doob’s nonnegative supermartingale inequality, similar to the one stated in Theorem 16 of [3] for discrete-time systems, that have been developed in previous studies to examine upper bounds of the reachability probabilities. Moreover, it is not possible to adapt the constraint $\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]-v(\bm{x})\leq-\delta,\forall\bm{x}\in\mathcal{X}$ from equation (10) to $\mathcal{A}v(\bm{x})\leq-\delta,\forall\bm{x}\in\mathcal{X}$ in order to compute upper bounds for the liveness probabilities. If we consider that $\mathcal{A}v(\bm{x})\leq-\delta,\forall\bm{x}\in\mathcal{X}$ is true, then $\widehat{\mathcal{A}}v(\bm{x})\leq-\delta,\forall\bm{x}\in\mathcal{X}$ will also hold, which implies $\widehat{\mathcal{A}}v(\bm{x})\leq-\delta,\forall\bm{x}\in\overline{\mathcal{X}}$. However, this contradicts the condition $\widehat{\mathcal{A}}v(\bm{x})=0,\forall\bm{x}\in\partial\mathcal{X}$.
@@ 6.2 Equations Relaxation Based Invariance Verification
In this subsection, we present the second set of optimizations for addressing the safe probabilistic invariance verification problem in Definition 5.3. We begin by introducing an equation that characterizes the liveness probability of satying the safe set $\mathcal{X}$ for all the time. This equation is adapted from (20), where the unsafe set $\partial\mathcal{X}$ is regarded as the target set $\mathcal{X}_{r}$ in (20). We then propose two sufficient conditions for certifying lower and upper bounds of the liveness probabilities by relaxing the derived equation.
Lemma 6.2.
@@
(27) $\begin{cases}\widehat{\mathcal{A}}v(\bm{x})=0\\ v(\bm{x})=1_{\partial\mathcal{X}}(\bm{x})+\widehat{\mathcal{A}}u(\bm{x})\end{cases},$
@@
Thereby, $\mathbb{P}\big{(}\forall t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\big{)}=1-v(\bm{x}_{0})$, $\forall\bm{x}_{0}\in\mathcal{X}$.
Proof 6.3.
@@
Similar to the Section 4, two sufficient conditions can be obtained for certifying lower and upper bounds of the liveness probabilities via directly relaxing equation (27), respectively.
Theorem 6.4.
(28) $\begin{cases}v(\bm{x})\leq 1-\epsilon_{1}&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\leq 0&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq\mathcal{A}u(\bm{x})&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\geq 1&\forall\bm{x}\in\partial{\mathcal{X}}\end{cases},$
then $\mathbb{P}\big{(}\exists t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\partial\mathcal{X}\big{)}\leq v(\bm{x}_{0})\leq 1-\epsilon_{1}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$. Hence, $\mathbb{P}\big{(}\forall t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\big{)}\geq\epsilon_{1}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$.
Proof 6.5.
Similarly, we can show that constraints (26) and (28) are equivalent.
@@
Constraints (26) and (28) are equivalent.
Proof 6.7.
The proof is similar to Proposition 4.17, which is shown in Appendix A.
According to Theorem 6.4, a lower bound of the liveness probabilities can be computed via solving Op6 (shown in Appendix B).
The sufficient condition for certifying upper bounds of the liveness probabilities is presented in Theorem 6.8.
Theorem 6.8.
(29) $\begin{cases}v(\bm{x})\geq 1-\epsilon_{2}&\forall\bm{x}\in\mathcal{X}_{0}\\ \mathcal{A}v(\bm{x})\geq 0&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq\mathcal{A}u(\bm{x})&\forall\bm{x}\in\mathcal{X}\\ v(\bm{x})\leq 1&\forall\bm{x}\in\partial{\mathcal{X}}\end{cases},$
then $\mathbb{P}\big{(}\exists t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\partial\mathcal{X}\big{)}\geq v(\bm{x}_{0})\geq 1-\epsilon_{2}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$. Hence, $\mathbb{P}\big{(}\forall t\in\mathbb{R}_{\geq 0}.\bm{X}^{\bm{x}_{0}}(t,\bm{w})\in\mathcal{X}\big{)}\leq\epsilon_{2}$, $\forall\bm{x}_{0}\in\mathcal{X}_{0}$.
Proof 6.9.
@@
According to Theorem 6.8, an upper bound of the liveness probabilities can be computed via solving Op7 (shown in Appendix B).
@@ 7 Examples
In this section we demonstrate our theoretical developments on five examples. Since directly solving the problems Op0-Op8 is challenging, we relax them into semi-definite programs (SDP0-SDP8, see Appendix C) using the sum of squares decomposition for multivariate polynomials, which allows for efficient solving. We solve the resulting semi-definite programs using the tool Mosek 10.1.21 [4]. Furthermore, to ensure numerical stability during the solution of the semi-definite programs, we impose a constraint on the coefficients of the unknown polynomials ($v(\bm{x}),w(\bm{x}),u(\bm{x}),p(\bm{x}),s_{i}(\bm{x}),i=0,\ldots,4$) in SDP0-SDP8 and restrict them to the interval $[-100,100]$.
In addition, in the experiments we employ the Monte Carlo method to approximate the upper and lower bounds of the liveness probabilities. Specifically, we take a large number of samples (e.g., $10^{4}$) from the initial set and simulate the liveness probabilities of these samples over an extended time. The maximum probability among these samples is considered as the upper bound, while the minimum probability is considered as the lower bound. By using these bounds as ground truth values, we can assess the accuracy of the bounds computed via solving SDP0-SDP8.
d 2 6 10 14 18 22 26
$\epsilon_{1}$ by Op3 0.3574 0.6678 0.6917 0.7368 0.7575 0.7622 0.7647
$\epsilon_{2}$ by Op4 1.0000 0.9505 0.9488 0.9242 0.8991 0.8927 0.8771
$\epsilon_{1}$ by Op1 0.3574 0.6678 0.6917 0.7368 0.7575 0.7622 0.7647
$\epsilon_{1}$ by Op0 0.3574 0.5895 0.5937 0.6740 0.6867 0.7007 0.7284